Circular No. 7860

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002ap IN M74 G. H. Marion, P. Hoeflich, and J. C. Wheeler, University of Texas at

Austin; and W. Vacca, Max-Planck-Institut für Extraterrestrische Physik, report on near-infrared spectroscopy of SN 2002ap: “The spectrum (range 0.80–2.5 μm, resolution 1200) was obtained on Mar. 8.3 UT using the NASA Infrared Telescope Facility (+ SpeX). We see no evidence for Paschen lines or Brackett γ. We confirm the presence of a P-Cyg feature with a minimum at 1.03 μm (IAUC 7834, 7838), which could be attributed to He I (1.083 μm) or Mg II (1.0914, 1.0915, and 1.0950 μm). However, the He identification is not favored because we find no detection of He I at 2.058 μm, while there is a strong presence of other Mg features in the spectrum. Mg II is found at 2.13 and 1.58 μm, in blends with Si II. Mg I is prominent at 1.50 μm, and the wide absorption trough at 0.93-1.03 μm is plausibly a combination of Mg I, C I, and Mg II. Ca II is identified in P-Cyg features with absorption minima at 0.81 μm and (possibly blended with Mg I) at 1.16 μm. We also detect emission or P-Cyg features that are identified with C I at 0.91 μm, Si II at 1.29 μm, O I at 1.11 μm, and O I blended with Si II at 1.22 μm.”

COMET C/2002 E2 (SNYDER-MURAKAMI) Ephemeris extension to IAUC 7852 (elements MPC 2002-F56):

2002 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Mar. 27 19^{h}10.^{m}88 +15^{o}57.^{′}3 1.404 1.540 77.^{o}7 39.^{o}3 10.1
Apr. 1 19 15.14 +21 59.8 1.380 1.563 80.4 39.1 10.1
6 19 18.89 +28 11.5 1.368 1.588 82.6 38.7 10.2
11 19 22.02 +34 25.1 1.369 1.616 84.4 38.1 10.3
16 19 24.41 +40 32.9 1.383 1.646 85.7 37.4 10.4
21 19 25.85 +46 27.5 1.410 1.678 86.4 36.7 10.5
26 19 26.12 +52 02.7 1.449 1.713 86.5 35.9 10.6
May 1 19 24.88 +57 14.3 1.498 1.749 86.2 35.1 10.8
6 19 21.67 +61 59.4 1.555 1.787 85.5 34.3 11.0
11 19 15.81 +66 16.6 1.621 1.826 84.5 33.4 11.2
16 19 06.4 +70 05.0 1.692 1.867 83.3 32.5 11.4

Visual m_{1} estimates: Mar. 15.10 UT, 10.3 (B. H. Granslo, Eina, Norway, 0.25-m reflector); 17.09, 10.1 (A. Baransky, Kiev, Ukraine, 0.36-m reflector); 21.86, 10.9 (A. Pearce, Nedlands, W. Australia, 0.41-m reflector); 23.11, 9.9 (K. Hornoch, Lelekovice, Czech Republic, 0.13-m reflector).

2002 March 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green